# PVLM: Parsing-Aware Vision Language Model with Dynamic Contrastive Learning for Zero-Shot Deepfake Attribution

### v2

@@ IV-C Ablation Study
Effects of various modules. To evaluate the contribution of each module to learning ability, we observe the DFA performance on both seen and unseen generators using protocol 1. In Table VII, NE improves the performance of the attributor by 2.34% ACC on images created by unseen generators, showing that SRM noises provide significant clues to facilitate ZS-DFA. The improvement from incorporating EE (+2.48%) is clear, highlighting the importance of edge global embeddings. LE further boosts performance by about 21% ACC, when EE or PE is also involved. The gain brought by LE alone is marginal without the integration of EE or PE, as the effectiveness of LE relies on the synergy between the different features. LE provides fine-grained semantic guidance, which is enhanced when combined with the distinctive facial parsing prior information from PE or edge features from EE. Without these modules, LE’s potential is not fully realized, leading to a smaller performance gain. Besides, when the MPVE module (i.e. AE+NE+EE) is involved, the ACC of the model is 43.67% on unseen generators. The improvement from adding PE (+15.36%) is evident, highlighting the importance of discriminative facial parsing guidance. LE enhances performance by about 7.4% ACC on unseen generators. The combined gain from integrating both PE and LE (+22.76%) is particularly striking. We argue that fine-grained language embeddings from LE and facial attribute information from PE reinforce and enhance each other, synergistically guiding the model to capture both general and distinct attribution embeddings.
TABLE V: Effects of different contrastive center losses.
@@
TABLE VI: Effects of various face parsers.
@@
TABLE VII: Model ablation. ACC scores (%) on Protocol-1.
@@
Impacts of the DFACC loss. Models are trained with the VCC loss and the DFACC loss, respectively. In Table VI, the ACC of both CLIP and PVLM supervised by the VCC loss are 79.04% and 96.90% on protocol-1, respectively. By contrast, they achieve 89.07% ACC and 100% ACC under the constraints of our DFACC criteria, accordingly, which verifies the necessity of conducting intra-generator compactness and inter-generator separability based on the correlation among generators. We visualize features extracted from samples produced by various generators using t-SNE [41]. We randomly select 500 samples for each seen generator from the testing set of protocol-1. In Fig. 6, samples of each generator are clustered in respective domains. With the addition of the DFACC loss, samples within the generator are more compact, with those of highly correlated generators aggregated and those of weakly related ones separated. This shows that our DFACC loss realizes flexible intra-generator aggregation and inter-generator separation.
@@
Influences of the face parser. We delve into the impact of various face parsers, such as SegFace [42] and BiSeNet [10]. We test frameworks on protocol-1. In Table VI, the ACC reaches the maximum as the BiSeNet is involved. We argue that face parsing images extracted by BiSeNet are more discriminative than those encoded by other face parsers, which could boost DFA. We visualize parsing images generated by various face parsers, such as BiSeNet and SegFace, for fake faces produced by various generators. In Fig. 4, compared to SegFace, the BiSeNet face parser does not generalize well to unseen diffusion faces. There are more evident distinctions among various generator parsing images produced by BiSeNet, visually. We attribute this to BiSeNet’s design as a lightweight face parsing network, optimized for real-world data. Unlike more complex models, such as GAN-based architectures or vision transformers like SegFace, BiSeNet may lack the capacity and flexibility to handle the unique characteristics of generative models and may also struggle with the high-frequency details and artifacts typical of diffusion models.
@@
We further calculate the FID score of the various generator parsing images created by different face parsers. To ensure a fair and balanced evaluation, we maintain an equal distribution of genuine and forged face parsing images, at a 1:1 ratio. We use 10,000 real images and 10,000 fake images to calculate the FID score for each generator. Lower FID scores denote better image quality. Results across various generators are shown in Fig. 4. The FID score of diffusion parsing images from BiSeNet is higher than that of those from SegFace, showing that BiSeNet struggles to generalize to unseen diffusion faces. Despite the performance degradation of BiSeNet, in Table VI, the ACC score of the model with BiSeNet is higher than that of the model with SegFace. We argue that there are evident distinctions among various generator parsing distributions produced by BiSeNet, which can facilitate DFA.
@@
Impacts of different losses. To delve into the contribution of loss functions, we conduct ablations on different losses. The ablation result of various losses is displayed in Table X. When PVLM is supervised with purely DFA loss, the ACC is 35.76% on unseen generators, but an increase of about 4.4% in ACC could be achieved via adding the DFACC criterion. We argue that it could boost the generalizable deepfake attribution representations via flexible contrastive learning across various generators. At the same time, owing to the addition of the CMC loss, PVLM is improved by 20.05% ACC on unseen generators, demonstrating the effectiveness of language guidance. The introduction of the DCPC loss further increases the performance (+5.5%), which shows that the global face parsing features could be valuable prior information for both deepfake detection and attribution. Moreover, due to the integration of the KL loss, our PVLM model achieves an enhancement of 0.7% ACC, which indicates that it is valuable to conduct feature alignment between visual and textual modalities. The incorporation of the five criteria yields the best performance among these losses, indicating that the proposed loss realizes promising results.
@@
Effects of the margin $m$ in DFACC loss. In the DFACC loss, we measure the correlation between generators based on the margin value. We investigate the influence of different margin values on model performance. As shown in Table XII, the ACC score tends to improve with an increase of the margin value, indicating that a larger margin enhances the model’s ability to distinguish between generators, effectively. Notably, the performance reaches the peak when the margin value is set to 0.7. This suggests that a margin of 0.7 can serve as an optimal value for evaluating the correlation among generators.
Influences of the language encoder. To delve into the impact of LE, we conduct ablation experiments using protocol-1. The ablation results are shown in Table IX. The gains from introducing the LE module (+7.4%) are obvious. We argue that LE offers fine-grained and accurate language guidance via vision-language matching, which could enhance the DFA performance of models.
Impacts of the MMI module. To study the effect of MMI, we conduct ablation experiments using protocol-1. The ablation results are shown in Table XI. The gains from adding the MMI module (+3.26%) are evident. We argue that MMI facilitates diverse and comprehensive multi-view interactions, promoting the model to extract rich attribution features.
Effects of the fine-grained text prompts. To study the impact of fine-grained text prompts on model performance, we introduce learnable prompts with insufficient accuracy, to train PVLM and CLIP, respectively. In Table IX, the performance of both CLIP and PVLM declines significantly (-25.52% and -17.42%), due to the addition of learnable text prompts, accordingly. We argue that elaborate and accurate semantic guidance plays a vital role in fine-grained DFA tasks.
TABLE VIII: Effects of text prompts. LP is the learnable prompt.
@@
TABLE IX: Effects of the language encoder.
@@
TABLE X: Effects of various losses.
@@
TABLE XI: Effects of the MMI module.
@@
Influences of the factor $\lambda$ in DFACC loss. In the DFACC loss, we define the factor $\lambda$ to balance the loss term between $\mathcal{L}_{\mathrm{intra}}$ and $\mathcal{L}_{\mathrm{inter}}$. We study the effect of various factors on model performance. As Table XIII displays, the ACC score rises with the increase of factors, which indicates the effectiveness of flexible metric learning. Specifically, the performance reaches its peak when the weight factor $\lambda$ is set to 0.5, which means an optimal balance between the two loss terms. We argue that the optimal setting of $\lambda$ allows the model to learn the most discriminative features, leading to better generalization and more accurate predictions.
TABLE XII: Influence of margin values in DFACC loss.
@@
TABLE XIII: Influence of the factor in DFACC loss.
@@ VI Deepfake Detection
To study the significance of DFA, we evaluate the performance of the deepfake detectors and the deepfake attributors model using Protocol-2, respectively. The results of models on two tasks (i.e., deepfake detection and deepfake attribution) are displayed in Table XIV. It is noticed that the classification tends to become easier, with the number of categories increasing from 2 (deepfake detection) to 11 (deepfake attribution). We argue that there is a growing distinction between pristine and forged images, as the number of categories rises, which facilitates deepfake detection. For example, MFCLIP and DNA-Det obtain 99.74% ACC and 96.18% ACC on seen generators under the deepfake detection setting, respectively. By contrast, they achieve 99.91% ACC (+0.17%) and 99.65% ACC (+2.47%) on seen generators in DFA tasks, accordingly.
TABLE XIV: ACC results (%) of models for deepfake detection and deepfake attribution on protocol-2.
@@ VII Conclusion
In this paper, we conduct a novel fine-grained ZS-DFA benchmark. Besides, we propose an innovative parsing-aware vision language model with dynamic contrastive learning (PVLM) method, which introduces face parsing modality to enhance the learning of visual forgery features across image, noise, and edge views, thus mining both general and discriminative attribution patterns, to achieve ZS-DFA. Although our method has achieved excellent ZS-DFA, the interaction among vision, parsing, and language modalities is limited. We intend to explore diverse and efficient vision-language-parsing communication mechanisms to improve the ZS-DFA performance of PVLM. We also plan to introduce abundant face data generated by various generators like flow-based and auto-regressive models, to facilitate ZS-DFA. We hope that our new benchmark and PVLM model could provide new insights for the research direction, such as deepfake attribution, deepfake detection, and vision language model.

### v3

@@ IV-C Ablation Study
Effects of various modules. To evaluate the contribution of each module to learning ability, we observe the DFA performance on both seen and unseen generators using protocol 1. In Table VIII, NE improves the performance of the attributor by 2.34% ACC on images created by unseen generators, showing that SRM noises provide significant clues to facilitate ZS-DFA. The improvement from incorporating EE (+2.48%) is clear, highlighting the importance of edge global embeddings. LE further boosts performance by about 21% ACC, when EE or PE is also involved. The gain brought by LE alone is marginal without the integration of EE or PE, as the effectiveness of LE relies on the synergy between the different features. LE provides fine-grained semantic guidance, which is enhanced when combined with the distinctive facial parsing prior information from PE or edge features from EE. Without these modules, LE’s potential is not fully realized, leading to a smaller performance gain. Besides, when the MPVE module (i.e. AE+NE+EE) is involved, the ACC of the model is 43.67% on unseen generators. The improvement from adding PE (+15.36%) is evident, highlighting the importance of discriminative facial parsing guidance. LE enhances performance by about 7.4% ACC on unseen generators. The combined gain from integrating both PE and LE (+22.76%) is particularly striking. We argue that fine-grained language embeddings from LE and facial attribute information from PE reinforce and enhance each other, synergistically guiding the model to capture both general and distinct attribution embeddings.
TABLE VI: Effects of different contrastive center losses.
@@
TABLE VII: Effects of various face parsers.
@@
TABLE VIII: Model ablation. ACC scores (%) on Protocol-1.
@@
Impacts of the DFACC loss. Models are trained with the VCC loss and the DFACC loss, respectively. In Table VII, the ACC of both CLIP and PVLM supervised by the VCC loss are 79.04% and 96.90% on protocol-1, respectively. By contrast, they achieve 89.07% ACC and 100% ACC under the constraints of our DFACC criteria, accordingly, which verifies the necessity of conducting intra-generator compactness and inter-generator separability based on the correlation among generators. We visualize features extracted from samples produced by various generators using t-SNE [35]. We randomly select 500 samples for each seen generator from the testing set of protocol-1. In Fig. 8, samples of each generator are clustered in respective domains. With the addition of the DFACC loss, samples within the generator are more compact, with those of highly correlated generators aggregated and those of weakly related ones separated. This shows that our DFACC loss realizes flexible intra-generator aggregation and inter-generator separation.
@@
Influences of the face parser. We delve into the impact of various face parsers, such as SegFace [23] and BiSeNet [38]. We test frameworks on protocol-1. In Table VII, the ACC reaches the maximum as the BiSeNet is involved. We argue that face parsing images extracted by BiSeNet are more discriminative than those encoded by other face parsers, which could boost DFA. We visualize parsing images generated by various face parsers, such as BiSeNet and SegFace, for fake faces produced by various generators. In Fig. 6, compared to SegFace, the BiSeNet face parser does not generalize well to unseen diffusion faces. There are more evident distinctions among various generator parsing images produced by BiSeNet, visually. We attribute this to BiSeNet’s design as a lightweight face parsing network, optimized for real-world data. Unlike more complex models, such as GAN-based architectures or vision transformers like SegFace, BiSeNet may lack the capacity and flexibility to handle the unique characteristics of generative models and may also struggle with the high-frequency details and artifacts typical of diffusion models.
@@
We further calculate the FID score of the various generator parsing images created by different face parsers. To ensure a fair and balanced evaluation, we maintain an equal distribution of genuine and forged face parsing images, at a 1:1 ratio. We use 10,000 real images and 10,000 fake images to calculate the FID score for each generator. Lower FID scores denote better image quality. Results across various generators are shown in Fig. 6. The FID score of diffusion parsing images from BiSeNet is higher than that of those from SegFace, showing that BiSeNet struggles to generalize to unseen diffusion faces. Despite the performance degradation of BiSeNet, in Table VII, the ACC score of the model with BiSeNet is higher than that of the model with SegFace. We argue that there are evident distinctions among various generator parsing distributions produced by BiSeNet, which can facilitate DFA.
@@
Impacts of different losses. To delve into the contribution of loss functions, we conduct ablations on different losses. The ablation result of various losses is displayed in Table XI. When PVLM is supervised with purely DFA loss, the ACC is 35.76% on unseen generators, but an increase of about 4.4% in ACC could be achieved via adding the DFACC criterion. We argue that it could boost the generalizable deepfake attribution representations via flexible contrastive learning across various generators. At the same time, owing to the addition of the CMC loss, PVLM is improved by 20.05% ACC on unseen generators, demonstrating the effectiveness of language guidance. The introduction of the DCPC loss further increases the performance (+5.5%), which shows that the global face parsing features could be valuable prior information for both deepfake detection and attribution. Moreover, due to the integration of the KL loss, our PVLM model achieves an enhancement of 0.7% ACC, which indicates that it is valuable to conduct feature alignment between visual and textual modalities. The incorporation of the five criteria yields the best performance among these losses, indicating that the proposed loss realizes promising results.
@@
Effects of the margin $m$ in DFACC loss. In the DFACC loss, we measure the correlation between generators based on the margin value. We investigate the influence of different margin values on model performance. As shown in Table XIII, the ACC score tends to improve with an increase of the margin value, indicating that a larger margin enhances the model’s ability to distinguish between generators, effectively. Notably, the performance reaches the peak when the margin value is set to 0.7. This suggests that a margin of 0.7 can serve as an optimal value for evaluating the correlation among generators.
Influences of the language encoder. To delve into the impact of LE, we conduct ablation experiments using protocol-1. The ablation results are shown in Table X. The gains from introducing the LE module (+7.4%) are obvious. We argue that LE offers fine-grained and accurate language guidance via vision-language matching, which could enhance the DFA performance of models.
Impacts of the MMI module. To study the effect of MMI, we conduct ablation experiments using protocol-1. The ablation results are shown in Table XII. The gains from adding the MMI module (+3.26%) are evident. We argue that MMI facilitates diverse and comprehensive multi-view interactions, promoting the model to extract rich attribution features.
Effects of the fine-grained text prompts. To study the impact of fine-grained text prompts on model performance, we introduce learnable prompts with insufficient accuracy, to train PVLM and CLIP, respectively. In Table X, the performance of both CLIP and PVLM declines significantly (-25.52% and -17.42%), due to the addition of learnable text prompts, accordingly. We argue that elaborate and accurate semantic guidance plays a vital role in fine-grained DFA tasks.
TABLE IX: Effects of text prompts. LP is the learnable prompt.
@@
TABLE X: Effects of the language encoder.
@@
TABLE XI: Effects of various losses.
@@
TABLE XII: Effects of the MMI module.
@@
Influences of the factor $\lambda$ in DFACC loss. In the DFACC loss, we define the factor $\lambda$ to balance the loss term between $\mathcal{L}_{\mathrm{intra}}$ and $\mathcal{L}_{\mathrm{inter}}$. We study the effect of various factors on model performance. As Table XIV displays, the ACC score rises with the increase of factors, which indicates the effectiveness of flexible metric learning. Specifically, the performance reaches its peak when the weight factor $\lambda$ is set to 0.5, which means an optimal balance between the two loss terms. We argue that the optimal setting of $\lambda$ allows the model to learn the most discriminative features, leading to better generalization and more accurate predictions.
TABLE XIII: Influence of margin values in DFACC loss.
@@
TABLE XIV: Influence of the factor in DFACC loss.
@@ VI Deepfake Detection
To study the significance of DFA, we evaluate the performance of the deepfake detectors and the deepfake attributors model using Protocol-2, respectively. The results of models on two tasks (i.e., deepfake detection and deepfake attribution) are displayed in Table XV. It is noticed that the classification tends to become easier, with the number of categories increasing from 2 (deepfake detection) to 11 (deepfake attribution). We argue that there is a growing distinction between pristine and forged images, as the number of categories rises, which facilitates deepfake detection. For example, MFCLIP and DNA-Det obtain 99.74% ACC and 96.18% ACC on seen generators under the deepfake detection setting, respectively. By contrast, they achieve 99.91% ACC (+0.17%) and 99.65% ACC (+2.47%) on seen generators in DFA tasks, accordingly.
TABLE XV: ACC results (%) of models for deepfake detection and deepfake attribution on protocol-2.
@@ VII Conclusion
In this paper, we conduct a novel fine-grained ZS-DFA benchmark. Besides, we propose an innovative parsing-aware vision-language model with dynamic contrastive learning (PVLM) method, which introduces face parsing modality to enhance the learning of visual forgery features across image, noise, and edge views, thus mining both general and discriminative attribution patterns, to achieve ZS-DFA. Although our method has achieved excellent ZS-DFA, the interaction among vision, parsing, and language modalities is limited. We intend to explore diverse and efficient vision-language-parsing communication mechanisms to improve the ZS-DFA performance of PVLM. We also plan to introduce abundant face data generated by various generators like flow-based and auto-regressive models, to facilitate ZS-DFA. We hope that our new benchmark and PVLM model could provide new insights for the research direction, such as deepfake attribution, deepfake detection, and vision-language model.